# Degeneration of families of projective hypersurfaces and Hodge conjecture

### v10

@@ 1 Introduction
The main result of this article is an inductive proof of the Hodge conjecture using the results of [4] and [5]. More precisely, assuming the Hodge conjecture for smooth projective varieties of dimension less or equal to $d-1$, the property of satisfying the Hodge conjecture for connected smooth projective varieties of dimension $d$ is a birational invariant (proposition 1). On the other hand, if $X$ is a connected smooth projective variety of dimension $d$, then $X$ is birational to a possibly singular hypersurface $Y_{0_{X}}\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ of degree $l_{X}:=deg(\mathbb{C}(X)/\mathbb{C}(x_{1},\cdots,x_{d}))$. Consider
$f:\mathcal{Y}^{\prime}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\mathbb{Q}}\subset\mathbb{P}(l_{X})$ the $\mathbb{Q}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture by [5] and by the induction hypothesis since $X$ is birational to $Y_{0_{X}}$. Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\mathbb{Q}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\mathbb{Q}$ such that
@@
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup S$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup S$ contain inductively the singular locus of the intersection with $\Delta$ and $S$.
@@
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}\geq 1)$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}=0$). We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$ where $\bar{S}^{\prime}\subset\Delta$ is a divisor. We then consider the family of algebraic varieties defined over $\mathbb{Q}$
@@
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective, in particular $E_{0}$ is birational to $Y_{0_{X}}$, hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
@@
$Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o}),$
whose de Rham part are morphism of filtered vector bundle over $S$ with Gauss-Manin connexion defined over $\mathbb{Q}$. Moreover, we have, see lemma 1, an isomorphism of variation of mixed Hodge structures over $S$
@@
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$\tilde{\lambda}:=Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q})).$
@@
using the fact that there exists a neighborhood $V_{S\cup S^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $S_{\mathbb{C}}\cup S^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\mathcal{E}^{\prime}}:\mathcal{E}^{\prime}_{\mathbb{C}}\hookrightarrow\mathcal{Y}_{\mathbb{C},V_{S\cup S^{\prime}}}$ retraction $r:\mathcal{Y}_{\mathbb{C},V_{S\cup S^{\prime}}}\to\mathcal{E}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
@@
$Sp_{\mathcal{E}/\tilde{f}_{X}}(q^{-1}(\mathbb{V}_{S}^{p}(\lambda)))=\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}),$ (1)
@@
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(\tilde{\lambda})\subset E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
@@
    $Sp_{\mathcal{E}/\tilde{f}_{X}}(\mathbb{V}_{S}^{p}(\tilde{\lambda}))=\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}),$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{0}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection, is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, where the second equality follows from [6] lemma 2.11. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(q^{-1}(\mathbb{V}_{S}^{p}(\lambda)))=q(\mathbb{V}_{S}^{p}(\tilde{\lambda}))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
@@
We recall the key point is that since $\mathbb{V}_{S}^{p}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$ is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugate are also components of the locus of Hodge classes, we have for each $\theta\in Aut(\mathbb{C},\mathbb{Q})$,
$\theta(\lambda)\in Gal(\bar{\mathbb{Q}},\mathbb{Q})(\mathbb{V}_{S}^{p}(\lambda)_{\theta})\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C},\theta}/S_{\mathbb{C},\theta}),$
@@ 2.4 The nearby cycle functor and the specialization map
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int f(O_{X},F),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S)).$
@@
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int_{p}\Gamma^{Hdg}_{X}(O_{\mathbb{P}^{N}\times S},F),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S)).$
@@
    If $f=f^{\prime}\circ\epsilon$ where $\epsilon:(X,E)\to(X,X_{S})$ is a desingularization, we have $Sp_{X_{S}/X^{\prime}}=Sp_{E/X}\circ\epsilon^{*}$.
@@
  • $S:=T\backslash T^{o}$ is a divisor, $E:=f^{-1}(S)=\cup_{i=0}^{r}E_{i}\subset X$ is a normal crossing divisor and that for $I\subset[0,\ldots r]$, $f_{0,I}:=f_{|E_{I}}:E_{I}:=\cap_{i\in I}E_{i}\to S$ are smooth. We have by definition $X^{o}=X\backslash E$.
@@
    If $H^{2p-1}(X_{t},\mathbb{Q})=0$ for $t\in T\backslash S$, then

### v11

@@ 1 Introduction
The main result of this article is an inductive proof of the Hodge conjecture using the results of [4] and [5]. More precisely, assuming the Hodge conjecture for smooth complex projective varieties of dimension less or equal to $d-1$, the property of satisfying the Hodge conjecture for connected smooth complex projective varieties of dimension $d$ is a birational invariant (proposition 1). Let $X$ be a connected smooth complex projective variety of dimension $d$. then $X$ is birational to a possibly singular hypersurface $Y_{0_{X}}\subset\mathbb{P}^{d+1}_{\mathbb{C}}$ of degree $l_{X}:=deg(\mathbb{C}(X)/\mathbb{C}(x_{1},\cdots,x_{d}))$. Consider
$f:\mathcal{Y}^{\prime}_{\mathbb{C}}=V(\tilde{f})\subset\mathbb{P}^{d+1}_{\mathbb{C}}\times\mathbb{P}(l_{X})\to\mathbb{P}(l_{X})$
the universal family of projective hypersurfaces of degree $l_{X}$, and $0_{X}\in\mathbb{P}(l_{X})$ the point corresponding to $Y_{0_{X}}$. Denote $\Delta\subset\mathbb{P}(l_{X})$ the discriminant locus parametrizing the hypersurfaces which are singular. Let $\bar{S}:=\bar{0_{X}}^{\mathbb{Q}}\subset\mathbb{P}(l_{X})$ the $\bar{\mathbb{Q}}$-Zariski closure of $0_{X}$ inside $\mathbb{P}(l_{X})$. If $Y_{0_{X}}$ is smooth, then $X$ satisfy the hodge conjecture since $Y_{0_{X}}$ satisfy the Hodge conjecture by [5] and $X$ is birational to $Y_{0_{X}}$, using the induction hypothesis. Hence, we may assume that $Y_{0_{X}}$ is singular, that is $0_{X}\in\Delta$. Then, $\bar{S}\subset\Delta$ since $\Delta\subset\mathbb{P}(l_{X})$ is defined over $\mathbb{Q}$. Let $\bar{T}\subset\mathbb{P}(l_{X})$ be an irreducible closed subvariety of dimension $\mathop{\rm dim}\nolimits(\bar{S})+1$ defined over $\bar{\mathbb{Q}}$ such that
@@
We take for $\bar{T}=V(f_{1},\cdots,f_{s})\subset\mathbb{P}(l_{X})$ an irreducible complete intersection defined over $\bar{\mathbb{Q}}$ which intersect $\Delta$ properly and such that
  • $V(f_{1})\supset\mathop{\rm sing}\nolimits(\Delta)\cup\bar{S}$
  • for each $r\in[1,\cdots,s]$, $V(f_{1},\cdots,f_{r})\supset\mathop{\rm sing}\nolimits(V(f_{1},\cdots,f_{r-1})\cap\Delta)\cup\bar{S}$ contain inductively the singular locus of the intersection with $\Delta$ and $\bar{S}$.
@@
is an isomorphism if $\mathop{\rm dim}\nolimits\bar{T}\geq 2$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}\geq 1)$ and surjective if $\mathop{\rm dim}\nolimits\bar{T}=1$ (i.e. $\mathop{\rm dim}\nolimits\bar{S}=0$). We have then $\bar{T}\cap\Delta=\bar{S}\cup\bar{S}^{\prime}$, where $\bar{S}^{\prime}\subset\bar{T}$ is a divisor. We then consider the family of algebraic varieties defined over $\bar{\mathbb{Q}}$
@@
  • $t\in S(\mathbb{C})$ is the point such that $\mathcal{Y}^{\prime}_{\mathbb{C},t}=Y_{0_{X}}$ and $\epsilon_{t}:E_{0}:=\mathcal{E}_{0,\mathbb{C},t}\to Y_{0_{X}}$ is surjective (there is at least one component of $E$ dominant over $Y_{O_{X}}$ since $E$ is dominant over $Y_{O_{X}}$ as $\mathcal{E}$ is dominant over $\mathcal{Y}^{\prime}_{\bar{S}}$), in particular $E_{0}$ is birational to $Y_{0_{X}}$, hence $E_{0}$ is birational to $X$. We denote $E_{i}:=\mathcal{E}_{i,\mathbb{C},t}$ for each $0\leq i\leq r$ and $E:=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$.
@@
$Sp_{\mathcal{E}/\tilde{f}_{X}}:E^{2p}_{Hdg}(\mathcal{E}/S)\to H^{2p}\psi_{S}E_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o})=\psi_{S}E^{2p}_{Hdg}(\mathcal{Y}_{T^{o}}/T^{o}),$
whose de Rham part are morphism of filtered vector bundle over $S$ with Gauss-Manin connexion defined over $\bar{\mathbb{Q}}$. Moreover, we have, see lemma 1, an isomorphism of variation of mixed Hodge structures over $S$
@@
where $HL^{2p,p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$, $HL^{2p,p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})\subset E_{DR}^{2p}(\mathcal{E}_{i,\mathbb{C}}/S_{\mathbb{C}})$ are the locus of Hodge classes. Let $\lambda\in F^{p}H^{2p}(E_{0}^{an},\mathbb{Q})$, where we recall $E_{0}=\mathcal{E}_{0,\mathbb{C},t}$ and $E=\mathcal{E}_{\mathbb{C},t}=\cup_{i=0}^{r}E_{i}$. Consider then $\tilde{\lambda}\in H^{2p}(E^{an},\mathbb{Q})$, such that $q(\tilde{\lambda})=\lambda$, and
$Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})\in Sp_{E/f_{X}}(H^{2p}(E^{an},\mathbb{Q}))\subset i_{t}^{*}\psi_{S}R^{2p}\tilde{f}_{X*}\mathbb{Q}_{\mathcal{Y}_{T^{o},\mathbb{C}}^{an}}.$
@@
using the fact that there exists a neighborhood $V_{\bar{S}\cup\bar{S}^{\prime}}\subset\bar{T}_{\mathbb{C}}$ of $\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ in $\bar{T}_{\mathbb{C}}$ for the usual complex topology such that the inclusion $i_{\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}}:\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}\hookrightarrow V_{\bar{S}\cup\bar{S}^{\prime}}$ admits a retraction $r:V_{\bar{S}\cup\bar{S}^{\prime}}\to\bar{S}_{\mathbb{C}}\cup\bar{S}^{\prime}_{\mathbb{C}}$ which is an homotopy equivalence. The locus of Hodge classes passing through $\lambda$
@@
$\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda})))),$ (1)
@@
  • $\mathbb{V}_{S}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0,\mathbb{C}}/S_{\mathbb{C}})$, $\mathbb{V}_{S}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$, are the flat leaves, e.g. $\mathbb{V}_{S}(\lambda):=\pi_{S}(\lambda\times\tilde{S}^{an}_{\mathbb{C}})$ where $\pi_{S}:H^{2p}(E^{an},\mathbb{C})\times\tilde{S}_{\mathbb{C}}^{an}\to E^{2p,an}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})$ is the morphism induced by the universal covering $\pi_{S}:\tilde{S}_{\mathbb{C}}^{an}\to S_{\mathbb{C}}^{an}$,
@@
  • $Sp_{\mathcal{E}/\tilde{f}_{X}}:=Sp_{\mathcal{E}/\tilde{f}_{X}}\otimes\mathbb{C}:E^{2p}_{DR}(\mathcal{E}_{\mathbb{C}}/S_{\mathbb{C}})\to i_{S}^{*mod}E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o})$.
@@
    $\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))=\overline{\mathbb{V}_{T^{o}}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))}\cap p^{-1}(S_{\mathbb{C}})\subset E^{2p}_{DR,V_{S}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T_{\mathbb{C}}^{o}),$
    where $p:=p\otimes\mathbb{C}:E^{2p}_{DR,V_{0}}(\mathcal{Y}_{T^{o},\mathbb{C}}/T^{o}_{\mathbb{C}})\to\bar{T}^{o}_{\mathbb{C}}$ is the projection, is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugates are also components of the locus of Hodge classes, where equality follows from [6] lemma 2.11. Hence,
    $\mathbb{V}_{S}^{p}(\lambda)=q(Sp_{\mathcal{E}/\tilde{f}_{X}}^{-1}(\mathbb{V}_{S}^{p}(Sp_{\mathcal{E}/\tilde{f}_{X}}(\tilde{\lambda}))))\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$
@@
We recall the key point is that since $\mathbb{V}_{S}^{p}(\lambda)\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C}}/S_{\mathbb{C}})$ is an algebraic subvariety defined over $\bar{\mathbb{Q}}$ and its Galois conjugate are also components of the locus of Hodge classes, we have for each $\theta\in Aut(\mathbb{C}/\mathbb{Q})$,
$\theta(\lambda)\in Gal(\bar{\mathbb{Q}}/\mathbb{Q})(\mathbb{V}_{S}^{p}(\lambda)_{\theta})\subset E^{2p}_{DR}(\mathcal{E}_{0\mathbb{C},\theta}/S_{\mathbb{C},\theta}),$
@@ 2.4 The nearby cycle functor and the specialization map
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int_{f}(O_{X},F_{b}),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S))$
and $E^{j}_{Hdg}(X/S):=H^{j}E_{Hdg}(X/S)$ for $j\in\mathbb{Z}$. If moreover $Z\subset X$ is a closed subset, we denote
$E_{Hdg,Z}(X/S):=f_{*}\Gamma_{Z}^{Hdg}\mathbb{Z}_{X}^{Hdg}:=(\int_{f}\Gamma_{Z}^{Hdg}(O_{X},F_{b}),Rf_{*}R\Gamma_{Z}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\Gamma_{Z}\alpha(X))\in D(MHM_{k,gm}(S))$
@@
$E_{Hdg}(X/S):=f_{*}\mathbb{Z}_{X}^{Hdg}:=(\int_{p}\Gamma^{\vee,Hdg}_{X}(O_{\mathbb{P}^{N}\times S},F),Rf_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},f_{*}\alpha(X))\in D(MHM_{k,gm}(S))$
and $E^{j}_{Hdg}(X/S):=H^{j}E_{Hdg}(X/S)$ for $j\in\mathbb{Z}$. In the particular case where $E_{Hdg}^{j}(X/S)$ is a variation of mixed Hodge structure, the first factor of
$E^{j}_{Hdg}(X/S):=(E^{j}_{DR}(X/S),R^{j}f_{*}\mathbb{Z}_{X^{an}_{\mathbb{C}}},H^{j}f_{*}\alpha(X))\in MHM_{k,gm}(S)$
is a vector bundle.
@@
    Recall the map $Sp_{E/X}$ is the defined as the factorization
    $\mathbb{Z}_{X}^{Hdg}\xrightarrow{\mathop{\rm ad}\nolimits(i_{E}^{*},i_{E*})(\mathbb{Z}_{X}^{Hdg}}\mathbb{Z}_{E}^{Hdg}\xrightarrow{Sp_{E/X}}\psi_{E}\mathbb{Z}^{Hdg}_{X^{o}}.$
@@
    If $f=f^{\prime}\circ\epsilon$ where $\epsilon:(X,E)\to(X,X_{S})$ is a desingularization, we have $Sp_{X_{S}/f^{\prime}}=Sp_{E/f}\circ\epsilon^{*}$.
@@
  • $S:=T\backslash T^{o}=V(s)\subset T$ is a divisor, $E:=f^{-1}(S)=\cup_{i=0}^{r}E_{i}\subset X$ is a normal crossing divisor and that for $I\subset[0,\ldots r]$, $f_{0,I}:=f_{|E_{I}}:E_{I}:=\cap_{i\in I}E_{i}\to S$ are smooth. We have by definition $X^{o}=X\backslash E$.
@@
    If $H^{2p-1}(X_{t},\mathbb{Q})=H^{2p-2}(X_{t},\mathbb{Q})=0$ for $t\in T\backslash S$, then
    $c_{E}:E^{k-1}_{Hdg,E}(X/T)\xrightarrow{c_{V}}\phi_{S}E^{k-1}_{Hdg}(X^{o}/T^{o})$
    $\xrightarrow{H^{k-1}(f\circ i_{E})_{*}c(\mathbb{Q}^{Hdg}_{E})}\ker(sp_{E/f}:E^{k}_{Hdg}(E/S)\to\psi_{S}E^{k}_{Hdg}(X^{o}/T^{o})).$